Projection Admin 3/26



To do this evening/ was completed: The fire alarm went off this morning. Juan and some folks were working on a few things which I believe set it off, or something in the kitchen, but either way, all theatres have been reset! Every corner of this establishment has had quite the bacon-y smell today. - L

Presentation Issues:

Preshow Clips/Playlist: Staff screenings for DOAU and Eephus have been emailed out, scheduled on the LMS and blocked on the calendar. - Lillian

Built the playlists for Eephus, Discreet Charm, Minecraft (Hanna playlist can get built once it's done ingesting to the LMS. Thanks! - Lillian 

Trailer Updates: Ingested the trailer for How to Train Your Dragon, The Shrouds - Lil

Special Events/ AV set up: Made the 1.90 lens file in T1 -A

The private event for Superhuman Body on 4/4 (next week) has been. moved to 4/28, but the 4/03 screening stays! So, we'll need new keys for later in the month, but we'll have the current set all day on 4/3. Chenelle is aware. - Lillian 

Will show Lillian the slightly reorganized T1 under-curtain area tomorrow when we set up for the KGO run through -A

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): 
Ingested 12 Monkeys to the LMS. 
Downloaded and ingested the preshow for Session in Class
Downloaded and ingested the preshow for Minecraft (Video Games 25min/15min and Minecraft 5min)
Downloaded and ingested the trailer for The Shrouds
Ingested Hanna to the LMS/T2 
Ingested 12 Monkeys 
Ingested Return of the Living Dead
Received the print of Kings of Comedy, which is now in T1, ready to get inspected. - Lillian 

Outgoing/Shipping (content): Deleted the incomplete/unneeded features off of each server after the FilmVault gaffe. -A

Conquest print has been re-inspected and is outbound. - Lillian

Print/video Inspection: See above! - Lil

Asset tracking doc updated/checked?:

Maintenance/cleaning: Pretty much finished tidying cables etc. under the curtain in Theater 1. The need for this section of the floor to be mopped is frankly comical... will probably need to be hit multiple times before it's really clean. There should be a good amount of time to make the first pass after the last round tonight ;) -A

Put together the shelf behind the curtain in T2. The PA speakers and equipment bin are on it, but I still need to organize stuff back there -A

Cue sheets updated/checked (every shift): Added the formula information for Eephus and updated the sheets for this upcoming week of 3/28. Made some adjustments for tomorrow. - Lillian 

Marquee updated?:

Posters updated?: Received posters for Eephus and Freaky Tales, both of which have been labeled and stored in the cubbies in 3/7. - Lillian

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

Projection Admin (and Close) 10/19

NPP Projection Report Friday 1/31 CLOSE Chris